OpenStudy (anonymous):

y =  x + 4 –3x + 2y = 8 change into slope intercept form. help?

OpenStudy (anonymous):

do you know what slope int form is?

OpenStudy (anonymous):

no

OpenStudy (youngster):

it is y=mx+b

OpenStudy (youngster):

so just simplify on the left side.

OpenStudy (anonymous):

yes solve for y tell me what you get

OpenStudy (anonymous):

yeah can you show me how to do this

OpenStudy (youngster):

so the equation is y = x + 4–3x + 2y = 8 , right?

OpenStudy (youngster):

so simplify it to y= -2x+2y+4=8

OpenStudy (anonymous):

so -3x+2y=8 so need to get y by it self add 3x to both sides -3x+3x+2y=8+3x 2y=3x+8 now need y by it self so need to get rid of the 2 so divide 2/2 y= 3/2 x+8/2 y=3/2 x+4